Grade 10Organic Chemistry

CuSO4 + Fe = ?

vansh
11 Years agoGrade 10

4 Answers

KUNCHAM SAMPATH
11 Years ago
Fe2(So4)3(aq)CuSO4 + Fe =
KUNCHAM SAMPATH
11 Years ago
sorry.....CuSO4 + Fe = Fe2(So4)3(aq)
RAKESH CHINDAM
11 Years ago
CuSO4+Fe = Fe2(SO4)3(aq) 
SHANMUKESHWAR
11 Years ago
 CuSO4+Fe = Fe2(SO4)3(aq)